Characteristics of Production Process, Cost Measurement

Vince Melders, of EcoScape Company, designs and installs custom lawn and garden irrigation systems for homes and businesses throughout the state. Each job is different, requiring different materials and labor for installing the systems. EcoScape estimated the following for the year:

Number of direct labor hours 6,870
Direct labor cost $68,700
Overhead cost $49,464

During the year, the following actual amounts were experienced:

Number of direct labor hours 6,030
Direct labor incurred $66,330
Overhead incurred $48,240.00

Required:

1. Should EcoScape use process costing or job-order costing?

- Select your answer -Process costingJob-order costingItem 1

2. If EcoScape uses a normal costing system and overhead is applied on the basis of direct labor hours,

What is the overhead rate? Round your answer to the nearest cent.

$ per direct labor hour

What is the average actual wage rate?

$ per direct labor hour

What is the cost of an installation that takes $3,470 of direct materials and 20 direct labor hours?

$
Answer #1
Statement showing Computations
Particulars Amount
1) EcoScape should use job-order costing since Each job is different, requiring different materials and labor for installing the systems
2)
Overhead cost          49,464.00
Number of direct labor hours              6,870.00
overhead rate =49,464/6870                    7.20
Direct labor incurred          66,330.00
Number of direct labor hours              6,030.00
average actual wage rate =66330/6030                  11.00
Cost of installation:
direct materials            3,470.00
Direct Labour= 20*11                220.00
Overhead cost = 20*7.20                144.00
cost of an installation            3,834.00